# Higher property T and below-rank phenomena of lattices

### v1

@@ 3.2.2. Torsion growth
How does this connect to property T? Hutchcroft-Pete [hutchcroft+pete] show that every countable infinite property T group admits a free measure-preserving action with cost $1$. If Gaboriau’s fixed price conjecture would hold for property T groups, the cost would be independent of the specific free measure-preserving action. When compared with the projective limit action for a residually finite property T group $\Gamma$, we would obtain that $\lim_{i\to\infty}b_{1}(\Gamma;k)/[\Gamma:\Gamma_{i}]=0$. The fixed price conjecture remains open in the generality of property T groups.
Question 74.
@@
For higher-rank lattices we know more. Abért-Gelander-Nikolov [abert+gelander+nikolov] show that right-angled lattices in higher-rank simple Lie groups have vanishing rank and first $\mathbb{F}_{p}$-homology gradient. This includes all lattices of $\mathbb{Q}$}-rank at least $2$ but also some uniform lattices. Abért-Bergeron-Fraczyk-Gaboriau [abert+bergeron+fraczyk+gaboriau] show that the $\mathbb{F}_{p}$-homology and homology torsion gradients of a higher-rank lattice vanish in a range below the rational rank of $\Gamma$. Then Fraczyk-Mellick-Wilkens [fraczyk+mellick+wilkens] showed that all higher rank lattices satisfy Gaboriau’s fixed price conjecture and thus the expected vanishing rank and homology gradients in degree $1$ hold.
The following is the most far reaching conjecture that subsumes all the above, as well the idea of Benjamini-Schramm continuity of homology from [7samurai]. It was formulated in [li+loeh+moraschini+sauer+uschold]*Conjecture 1.15.
@@
Here are two wildly speculative attempts to tackle this conjecture. First, one might try to establish a fixed price conjecture for higher property T and the higher-dimensional cost, which is ongoing work of Abért-Gaboriau-Tanushevski, and a higher-dimensional analogue of the theorem of Pete-Hutchcroft. We do not expect that higher property T is sufficient but that specific facts about lattices are needed. However, one might speculate how far just higher property T carries. Second, one might try to establish a fixed price property for the measured embedding dimension and measured embedding volume defined by Li-Löh-Moraschini-Sauer-Uschold [li+loeh+moraschini+sauer+uschold] and use their bounds [li+loeh+moraschini+sauer+uschold]*Theorem 1.2 for the homology (torsion) gradients.
@@ 3.3. Geometric below-rank phenomena
3.3.1. Polynomiality of filling function below the rank
The proofs of Theorems 1 and 63 for non-uniform lattices have to solve the following problem. How to transfer cohomological information about the ambient Lie group $G$ to the lattice and back in the absence of a Shapiro lemma that generally is only available for uniform lattices?
The idea in [badsau] to circumvent this difficulty is to use polynomial cohomology as an auxiliary cohomology. Polynomial cohomology $H_{pol}^{\ast}(\Gamma,V)$ is defined using cochains in the bar complex that satisfy a polynomial growth constraint with respect to the word length. The comparison of polynomial and usual cohomology is via geometric group theory.
The $d$-th filling function $F_{d}^{\Gamma}(n)$ at $n>0$ of a group $\Gamma$ bounds the minimal (combinatorial) volume of a $d$-chain bounding a (d1)(d-1)-cycle of volume at most $n$ in a classifying space of finite type. The following is proved in [badsau]*Proposition 6.14.
@@ 3.3.1. Polynomiality of filling function below the rank
The following deep result by Leuzinger-Young [leuzinger+young] allows us to apply the previous theorem.
@@
Morally, a kind of Shapiro lemma for the polynomial cohomology with unitary or Banach space coefficients is proved in [badsau]. Via the two previous theorems, we obtain a suitable Shapiro lemma for non-uniform lattices below the rank.
In the S-arithmetic case and in positive characteristic, the equivalent of Theorem 77 is not available. Bestvina-Eskin-Wortman [bestvina+eskin+wortman] show polynomiality of non-uniform lattices in S-arithmetic groups not below the $S$-rank but up to a degree bounded by the number of primes in $S$. Further, Sauer-Weis [sauer+weis] show polynomiality below the rank (instead of the $S$-rank). Depending on the number of primes in $S$, one or the other result might be stronger.
@@
From the techniques in [badsau] and [BBBS] one can conclude the following.
@@
One could try to generalize Harder’s method in [harder] to prove Theorem 2 in positive characteristic. See also Remark 3. The following theorem shows an alternative route to Theorem 1 in positive characteristic, assuming Conjecture 78. We refer to our forthcoming work [BBBS] for details.
@@
Regarding the property T assumption, see the discussion in §3.4.1 below. Finally, we remark that in the case where the number of inverted primes in $S$ goes to infinity the range gap between the theorems of Leuzinger-Young and Bestvina-Eskin-Wortman becomes irrelevant. In the limiting, adelic case one can prove, for example, the following clean statement [badsau], which generalizes work of Borel and Yang.
@@ 3.3.2. Farb’s property FAn and speculations on the Zimmer program
A group satisfies Serre’s property FA if each of its actions on a tree has a fixed point. It is well known that property T groups satisfy Serre’s property FA. In [farb2009group] Farb makes the following generalization.
@@
Theorem 83 (Farb’s conjecture, [BBBS]).
@@
Theorems 83 and 84 fit Zimmer’s philosophy whose best-known representative is the following.
@@ 3.3.3. Waist inequalities below the rank
Waist inequalities predict the existence of a fiber with a uniform lower complexity bound for a family of maps to some Euclidean spaces, or more generally, manifolds. Gromov’s famous waist inequality [gromov-waist] for the sphere $S^{n}$ says that the maximal volume of a fiber of a (generic) map from $S^{n}$ to $d$-dimensional Euclidean space is at least the (nd)(n-d)-dimensional volume of an equator sphere $S^{n-d}$. The focus of the following discussion is less on explicit lower complexity bounds but allows families of maps where the domain manifold varies. A typical example is to consider all finite covers of a fixed manifold in the domain. In this setting waist inequalities can be understood as Riemannian versions of higher-dimensional topological expanders.
In this subsection we present a conjectural below-rank phenomenon for higher rank lattices and relate it to higher property T. A differential-geometric approach to waist inequalities was developed by Fraczyk and Lowe [fraczyklowe]. Our account is influenced by the connection to higher property T. See Figure 1 for a diagram illustrating much of the content of this subsection and the previous one.
@@
Using geometric measure theory and the theory of minimal surfaces, the following result was proved by Fraczyk and Lowe [fraczyklowe].
@@
With completely different methods, the authors prove the following result [BS-Waist].
Theorem 90.
@@
The proof of Theorem 90 relies on the concept of coboundary expansion with respect to cellular $\ell^{1}$-norm in the simplicial or cellular cochain complexes – a concept first studied by Linial-Meshulam. We call a CW-complex $X$ an $\mathbb{R}$}-valued $\epsilon$-coboundary expander in degree $k$ if for every cochain $g$ in the image of the differential $d\colon C^{k}(X;{\mathbb{R}})\to C^{k+1}(X;{\mathbb{R}})$ there is $f\in C^{k}(X;{\mathbb{R}})$ such that $d(f)=g$ and $\|g\|\geq\epsilon\cdot\|f\|$. Obvious variations yield the definitions of $\mathbb{Z}$}-valued or $\mathbb{F}_{2}$-valued $\epsilon$-coboundary expander. A $\mathbb{Z}$}-valued $\epsilon$-coboundary expander in degree $k$ such that each integral $k$-cocycle that is non-zero in cohomology has norm at least $\mu$ is called a $\mathbb{Z}$}-valued $(\epsilon,\mu)$-cosystolic expander. Compare with [evra+kaufman]*Definition 1.7.
For the proof of Theorem 90 one fixes a triangulation of the base manifold $M$ and endows every element of the family $\mathcal{F}$ consisting of the finite covers of $M$ with the lifted triangulations. The first step is to apply the following theorem to every triangulated manifold in $\mathcal{F}$ and for $k\in\{0,1\}$. The case $k\in\{0,1\}$ is proved in [BS-Waist] but there is a similar proof in arbitrary degrees.
@@
We show in [BS-Waist]*Theorem 2.12 for a compact CW-complex $X$ with $H^{1}(X,{\mathbb{Z}})=0$ that $\mathbb{R}$}-valued $\epsilon$-coboundary expansion implies $\mathbb{Z}$}-valued $\epsilon$-cosystolic expansion in degrees 0 and $1$. Together with Theorem 93, one concludes the proof of Theorem 90. Our tool for extracting $\mathbb{Z}$-coboundary expansion out of the $\mathbb{R}$-coboundary expansion given in Theorem 92 is the total unimodularity of certain coboundary maps, see [BS-Waist, Theorem 2.8]. Without overly expanding on this at this point, we remark on the relation between total unimodularity of the boundary maps and homological torsion, which was established in [THK, Theorem 5.2]. It is also implicit in the recent preprint [torsion_jakob]. We cannot help speculating that one might combine $(T_{n})_{L^{1}}$ and the sublinearity of torsion growth from Conjecture 75 to establish $\mathbb{Z}$ cohomological isoperimetric inequalities for lattices.
@@
Let us turn again to Conjectures 88 and 91. Using Corollary 8 we can apply Theorem 92 in both cases to get $\mathbb{R}$}-valued coboundary expansion. What is still missing is the transition to $\mathbb{Z}$} and from $\mathbb{Z}$}-valued coboundary expansion to $\mathbb{Z}$}-valued cosystolic expansion.
Question 96.
Under which conditions can one deduce $\mathbb{Z}$}-valued coboundary expansion from $\mathbb{R}$}-valued cosystolic expansion? Are there combinatorial-homological methods, reminiscent of Evra-Kaufman’s local-to-global theorem [evra+kaufman], that allow to further deduce $\mathbb{Z}$}-valued cosystolic expansion?
@@ 3.4.1. Degree 1: The spectral gap conjecture
Let $V$ be a unitary representation of $\Gamma$. If $V$ does not have a spectral gap, then $H^{1}(\Gamma,V)$ is non-Hausdorff and in particular non-zero. So in this case Conjecture 6 implies the existence of a non-trivial subrepresentation of $V$ on which the $\Gamma$-action extends to a $G$-action. For the reverse direction, assume that $H^{1}(\Gamma,V)\neq 0$ and that the $\Gamma$-action on $V$ does not extend to a $G$-action on any non-trivial subrepresentation. We now consider the unitary induction $I(V)$ of $V$ from $\Gamma$ to $G$. We have $H^{1}(G,I(V))\neq 0$. If $\Gamma<G$ is uniform, this follows from the Shapiro lemma. If $\Gamma<G$ is non-uniform, one can use the replacement of the Shapiro lemma mentioned in §3.3.1. By [BBHP, Proposition 5.1], we have that for every factor $G_{1}<G$, $I(V)^{G_{1}}=0$. We conclude by a result of Shalom [Shalom]*Proposition 3.2 that $H^{1}(G,I(V))$ is non-Hausdorff. Thus $I(V)$ does not have a spectral gap. Therefore, $V$ does not have a spectral gap. ∎
@@
Consider a non-compact factor $G_{1}<G$ with property T. Assume that $V$ does not have a spectral gap. Then its unitary induction $I(V)$ to $G$ does not have a spectral gap as $G$-representation. Hence also not as a $G_{1}$-representation. We deduce that $I(V)^{G_{1}}\neq 0$. However, by [BBHP, Proposition 5.1] $I(V)$ could be identified with a subrepresentation of $V$ on which the $\Gamma$-action extends to a $G$-action. ∎
@@
Clearly, Conjecture 98 implies Conjecture 97 and it holds if $G$ has property T. If $G$ is a Lie group, the argument of [badsau, Theorem D] shows the theorem still holds without assuming property T - provided Conjecture 98, see [badsau, Conjecture 1.5]. That is, we have:
@@
We list below various conjectures regarding irreducible higher rank lattices which are all implied by Conjecture 97 and in particular by Conjecture 6. For convenience, see Figure 2 for a summary of the relation between the various conjectures discussed in this subsection. We start with the following conjecture which is due to Lubotzky.
@@
The following conjecture is due to Cornulier. Some progress in this direction was recently made in [BGL].
@@
The following is a conjecture of Stuck-Zimmer [stuck+zimmer].
@@
In [BBHP] the concepts of charmenability and charfiniteness are defined. We will not repeat the definitions here. It is shown there that every irreducible higher rank lattice is charmenable and conjectured that it is charfinite.
@@
Before this we need some preparation, taken from [BSS].
Let $G$ be a semisimple group. For every unitary representation $V$ of $G$ we let $V_{\operatorname{temp}}$ be the unique maximal subrepresentation that is tempered (that is, weakly contained in the regular representation). We let $V_{\operatorname{rig}}$ be its orthogonal complement, thus $V=V_{\operatorname{temp}}\oplus V_{\operatorname{rig}}$. If $V=V_{\operatorname{rig}}$ we say that the representation $V$ is of rigid type. The following theorem explains why.
Theorem 104 ([BSS]).
@@
By the main results of [BBHP] and [BH19], $\Gamma$ is charmenable, thus it is charfinite if it satisfies items (3)-(5) in [BBHP, Definition 1.2]. Item (3) is clear by Zariski density and (4) follows from super-rigidity. So we only need to establish (5), that is, to show that every amenable extremal character of $\Gamma$ is finite. To this end, let us fix an amenable extremal character $\phi$.
@@
See [BBHP]. ∎
In case $G$ has a factor with T, Conjecture 103 was proved in [BBHP] and [BH19]. Very recently it was proved with no conditions on $G$ for all non-uniform lattices by Dogon-Glasner-Gorfine-Hanany-Levit [5authors]. However, it is still open for all cocompact lattices in products of groups which do not have T.
[…]

### v2

@@ 3.2.3. Torsion growth
How does this connect to property T? Hutchcroft-Pete [62] show that every countable infinite property T group admits a free measure-preserving action with cost $1$. If Gaboriau’s fixed price conjecture would hold for property T groups, the cost would be independent of the specific free measure-preserving action. When compared with the projective limit action for a residually finite property T group $\Gamma$, we would obtain that $\lim_{i\to\infty}b_{1}(\Gamma;k)/[\Gamma:\Gamma_{i}]=0$. The fixed price conjecture remains open in the generality of property T groups.
Question 82.
@@
For higher-rank lattices we know more. Abért-Gelander-Nikolov [3] show that right-angled lattices in higher-rank simple Lie groups have vanishing rank and first $\mathbb{F}_{p}$-homology gradient. This includes all lattices of $\mathbb{Q}$}-rank at least $2$ but also some uniform lattices. Abért-Bergeron-Fraczyk-Gaboriau [2] show that the $\mathbb{F}_{p}$-homology and homology torsion gradients of a higher-rank lattice vanish in a range below the rational rank of $\Gamma$. Then Fraczyk-Mellick-Wilkens [48] showed that all higher rank lattices satisfy Gaboriau’s fixed price conjecture and thus the expected vanishing rank and homology gradients in degree $1$ hold.
The following is the most far reaching conjecture that subsumes all the above, as well the idea of Benjamini-Schramm continuity of homology from [1]. It was formulated in [70]*Conjecture 1.15.
@@
Here are two wildly speculative attempts to tackle this conjecture. First, one might try to establish a fixed price conjecture for higher property T and the higher-dimensional cost, which is ongoing work of Abért-Gaboriau-Tanushevski, and a higher-dimensional analogue of the theorem of Pete-Hutchcroft. We do not expect that higher property T is sufficient but that specific facts about lattices are needed. However, one might speculate how far just higher property T carries. Second, one might try to establish a fixed price property for the measured embedding dimension and measured embedding volume defined by Li-Löh-Moraschini-Sauer-Uschold [70] and use their bounds [70]*Theorem 1.2 for the homology (torsion) gradients.
@@ 3.3. Geometric below-rank phenomena
3.3.1. Polynomial cohomology and polynomiality of filling function below the rank
The proofs of Theorems 1 and 69 for non-uniform lattices have to solve the following problem. How to transfer cohomological information about the ambient Lie group $G$ to the lattice and back in the absence of a Shapiro lemma that generally is only available for uniform lattices?
The idea in [15] to circumvent this difficulty is to use polynomial cohomology as an auxiliary cohomology. Polynomial cohomology $H_{pol}^{\ast}(\Gamma,V)$ is defined using cochains in the bar complex that satisfy a polynomial growth constraint with respect to the word length. The comparison of polynomial and usual cohomology is via geometric group theory.
The $d$-th filling function $F_{d}^{\Gamma}(n)$ at $n>0$ of a group $\Gamma$ bounds the minimal (combinatorial) volume of a $d$-chain bounding a (d1)(d-1)-cycle of volume at most $n$ in a classifying space of finite type. The following is proved in [15]*Proposition 6.14.
@@ 3.3.1. Polynomial cohomology and polynomiality of filling function below the rank
The following deep result by Leuzinger-Young [68] allows us to apply the previous theorem.
@@
Morally, a kind of Shapiro lemma for the polynomial cohomology with unitary or Banach space coefficients is proved in [15]. Via the two previous theorems, we obtain a suitable Shapiro lemma for non-uniform lattices below the rank.
In the S-arithmetic case and in positive characteristic, the equivalent of Theorem 85 is not available. Bestvina-Eskin-Wortman [18] show polynomiality of non-uniform lattices in S-arithmetic groups not below the $S$-rank but up to a degree bounded by the number of primes in $S$. Further, Sauer-Weis [90] show polynomiality below the rank (instead of the $S$-rank). Depending on the number of primes in $S$, one or the other result might be stronger.
@@
From the techniques in [15] and [7] one can conclude the following.
@@
One could try to generalize Harder’s method in [58] to prove Theorem 2 in positive characteristic. See also Remark 3. The following theorem shows an alternative route to Theorem 1 in positive characteristic, assuming Conjecture 86. We refer to our forthcoming work [7] for details.
@@
Regarding the property T assumption, see the discussion in §3.4.1 below. Finally, we remark that in the case where the number of inverted primes in $S$ goes to infinity the range gap between the theorems of Leuzinger-Young and Bestvina-Eskin-Wortman becomes irrelevant. In the limiting, adelic case one can prove, for example, the following clean statement [15], which generalizes work of Borel and Yang.
@@
Finally, we would like to mention a very recent development in the study of polynomial cohomology: López Neumann and Paucar [72] used polynomial cohomology and ideas in [15] to prove interesting new invariance results for Betti numbers of nilpotent groups in the setting of quantitative measure equivalence. Further, they develop techniques to deal with unitary induction for non-uniform lattices in Lie groups of rank $1$. See also the remark below Theorem 5.
@@ 3.3.2. Farb’s property FAn and speculations on the Zimmer program
A group satisfies Serre’s property FA if each of its actions on a tree has a fixed point. It is well known that property T groups satisfy Serre’s property FA. In [43] Farb makes the following generalization.
@@
Theorem 91 (Farb’s conjecture, [7]).
@@
Theorems 91 and 92 fit Zimmer’s philosophy whose best-known representative is the following.
@@ 3.3.3. Waist inequalities below the rank
Waist inequalities predict the existence of a fiber with a uniform lower complexity bound for a family of maps to some Euclidean spaces, or more generally, manifolds. Gromov’s famous waist inequality [56] for the sphere $S^{n}$ says that the maximal volume of a fiber of a (generic) map from $S^{n}$ to $d$-dimensional Euclidean space is at least the (nd)(n-d)-dimensional volume of an equator sphere $S^{n-d}$. The focus of the following discussion is less on explicit lower complexity bounds but allows families of maps where the domain manifold varies. A typical example is to consider all finite covers of a fixed manifold in the domain. In this setting waist inequalities can be understood as Riemannian versions of higher-dimensional topological expanders.
In this subsection we present a conjectural below-rank phenomenon for higher rank lattices and relate it to higher property T. A differential-geometric approach to waist inequalities was developed by Fraczyk and Lowe [47]. Our account is influenced by the connection to higher property T. See Figure 1 for a diagram illustrating much of the content of this subsection and the previous one.
@@
Using geometric measure theory and the theory of minimal surfaces, the following result was proved by Fraczyk and Lowe [47].
@@
With completely different methods, the authors prove the following result.
Theorem 98 ([16]).
@@
In a very recent preprint, Cameron Gates Rudd [89] uses ideas from [16] to prove that a residually finite $3$-dimensional Poincaré duality group cannot be Kähler.
@@
The proof of Theorem 98 relies on the concept of coboundary expansion with respect to cellular $\ell^{1}$-norm in the simplicial or cellular cochain complexes – a concept first studied by Linial-Meshulam. We call a CW-complex $X$ an $\mathbb{R}$}-valued $\epsilon$-coboundary expander in degree $k$ if for every cochain $g$ in the image of the differential $d\colon C^{k}(X;{\mathbb{R}})\to C^{k+1}(X;{\mathbb{R}})$ there is $f\in C^{k}(X;{\mathbb{R}})$ such that $d(f)=g$ and $\|g\|\geq\epsilon\cdot\|f\|$. Obvious variations yield the definitions of $\mathbb{Z}$}-valued or $\mathbb{F}_{2}$-valued $\epsilon$-coboundary expander. A $\mathbb{Z}$}-valued $\epsilon$-coboundary expander in degree $k$ such that each integral $k$-cocycle that is non-zero in cohomology has norm at least $\mu$ is called a $\mathbb{Z}$}-valued $(\epsilon,\mu)$-cosystolic expander. Compare with [41]*Definition 1.7.
For the proof of Theorem 98 one fixes a triangulation of the base manifold $M$ and endows every element of the family $\mathcal{F}$ consisting of the finite covers of $M$ with the lifted triangulations. The first step is to apply the following theorem to every triangulated manifold in $\mathcal{F}$ and for $k\in\{0,1\}$. The case $k\in\{0,1\}$ is proved in [16] but there is a similar proof in arbitrary degrees.
@@
We show in [16]*Theorem 2.12 for a compact CW-complex $X$ with $H^{1}(X,{\mathbb{Z}})=0$ that $\mathbb{R}$}-valued $\epsilon$-coboundary expansion implies $\mathbb{Z}$}-valued $\epsilon$-cosystolic expansion in degrees 0 and $1$. Together with Theorem 101, one concludes the proof of Theorem 98. Our tool for extracting $\mathbb{Z}$-coboundary expansion out of the $\mathbb{R}$-coboundary expansion given in Theorem 100 is the total unimodularity of certain coboundary maps, see [16, Theorem 2.8]. Without overly expanding on this at this point, we remark on the relation between total unimodularity of the boundary maps and homological torsion, which was established in [38, Theorem 5.2]. It is also implicit in the recent preprint [94]. We cannot help speculating that one might combine $(T_{n})_{L^{1}}$ and the sublinearity of torsion growth from Conjecture 83 to establish $\mathbb{Z}$ cohomological isoperimetric inequalities for lattices.
@@
Let us turn again to Conjectures 96 and 99. Using Corollary 12 we can apply Theorem 100 in both cases to get $\mathbb{R}$}-valued coboundary expansion. What is still missing is the transition to $\mathbb{Z}$} and from $\mathbb{Z}$}-valued coboundary expansion to $\mathbb{Z}$}-valued cosystolic expansion.
Question 104.
Under which conditions can one deduce $\mathbb{Z}$}-valued coboundary expansion from $\mathbb{R}$}-valued cosystolic expansion? Are there combinatorial-homological methods, reminiscent of Evra-Kaufman’s local-to-global theorem [41], that allow to further deduce $\mathbb{Z}$}-valued cosystolic expansion?
@@ 3.4.1. Degree 1: The spectral gap conjecture
Let $V$ be a unitary representation of $\Gamma$. If $V$ does not have a spectral gap, then $H^{1}(\Gamma,V)$ is non-Hausdorff and in particular non-zero. So in this case Conjecture 8 implies the existence of a non-trivial subrepresentation of $V$ on which the $\Gamma$-action extends to a $G$-action. For the reverse direction, assume that $H^{1}(\Gamma,V)\neq 0$ and that the $\Gamma$-action on $V$ does not extend to a $G$-action on any non-trivial subrepresentation. We now consider the unitary induction $I(V)$ of $V$ from $\Gamma$ to $G$. We have $H^{1}(G,I(V))\neq 0$. If $\Gamma<G$ is uniform, this follows from the Shapiro lemma. If $\Gamma<G$ is non-uniform, one can use the replacement of the Shapiro lemma mentioned in §3.3.1. By [9, Proposition 5.1], we have that for every factor $G_{1}<G$, $I(V)^{G_{1}}=0$. We conclude by a result of Shalom [91]*Proposition 3.2 that $H^{1}(G,I(V))$ is non-Hausdorff. Thus $I(V)$ does not have a spectral gap. Therefore, $V$ does not have a spectral gap. ∎
@@
Consider a non-compact factor $G_{1}<G$ with property T. Assume that $V$ does not have a spectral gap. Then its unitary induction $I(V)$ to $G$ does not have a spectral gap as $G$-representation. Hence also not as a $G_{1}$-representation. We deduce that $I(V)^{G_{1}}\neq 0$. However, by [9, Proposition 5.1] $I(V)$ could be identified with a subrepresentation of $V$ on which the $\Gamma$-action extends to a $G$-action. ∎
@@
Clearly, Conjecture 106 implies Conjecture 105 and it holds if $G$ has property T. If $G$ is a Lie group, the argument of [15, Theorem D] shows the theorem still holds without assuming property T - provided Conjecture 106, see [15, Conjecture 1.5]. That is, we have:
@@
We list below various conjectures regarding irreducible higher rank lattices which are all implied by Conjecture 105 and in particular by Conjecture 8. For convenience, see Figure 2 for a summary of the relation between the various conjectures discussed in this subsection.
[…]
We start with the following conjecture which is due to Lubotzky.
@@
The following conjecture is due to Cornulier. Some progress in this direction was recently made in [11].
@@
The following is a conjecture of Stuck-Zimmer [92].
@@
In [9] the concepts of charmenability and charfiniteness are defined. We will not repeat the definitions here. It is shown there that every irreducible higher rank lattice is charmenable and conjectured that it is charfinite.
@@
Before this we need some preparation, taken from [14]. Let $G$ be a semisimple group. For every unitary representation $V$ of $G$ we let $V_{\operatorname{temp}}$ be the unique maximal subrepresentation that is tempered (that is, weakly contained in the regular representation). We let $V_{\operatorname{rig}}$ be its orthogonal complement, thus $V=V_{\operatorname{temp}}\oplus V_{\operatorname{rig}}$. If $V=V_{\operatorname{rig}}$ we say that the representation $V$ is of rigid type. The following theorem explains why.
Theorem 112 ([14]).
@@
By the main results of [9] and [27], $\Gamma$ is charmenable, thus it is charfinite if it satisfies items (3)-(5) in [9, Definition 1.2]. Item (3) is clear by Zariski density and (4) follows from super-rigidity. So we only need to establish (5), that is, to show that every amenable extremal character of $\Gamma$ is finite. To this end, let us fix an amenable extremal character $\phi$.
@@
See [9]. ∎
In case $G$ has a factor with T, Conjecture 111 was proved in [9] and [27]. Very recently it was proved with no conditions on $G$ for all non-uniform lattices by Dogon-Glasner-Gorfine-Hanany-Levit [39]. However, it is still open for all cocompact lattices in products of groups which do not have T.